OpenStudy (anonymous):

this is rather very simple algebra question. #1) sam drives 20 miles at the rate of 60mph and then reduced the speed to 40mps at bumpy road for another 8 miles drive. What is the average speed of the entire driving?

OpenStudy (anonymous):

First calculate the total time taken for the entire diving. Total time taken = 20/60 + 8/40 = 8/15 hours

OpenStudy (anonymous):

Now, average speed = distance travelled / time taken =28/(8/15) =52.5mph

OpenStudy (anonymous):

Thank you so much!! I have got mixed up with something else here and went circling around. thanks for the help!1

OpenStudy (anonymous):

welcome